# Daniel Calegari and Andrea Delgado

### v1

@@ 3 Data synthesis & results
As already described and summarized in Table 3, the main objective addressed in the papers is the discovery of educational trajectories. It requires describing a process model that reflects the trajectory of students. Cases are usually actual students’ trajectories in which events of the discovered process represent courses, e.g., in 47, 2, 4, 13, 39, 9, 27, 19, or certain curriculum milestones such as the end of a semester 30. However, the semantics of an event can be conceived from different perspectives, not only from a curricular standpoint, e.g., the courses a student approves, but also from a contextual perspective, e.g., graduation in time, late dropouts, or the number of pending studies. This interpretation is used in works focusing on stopout/dropout analysis, e.g., 35, as we will describe later.
@@
Since a trajectory refers both to a student’s path and a set of students, in some cases, there is an a priori clustering of traces based on considering additional information, e.g., demographic information or students’ performance characteristics, which sometimes requires additional data mining tools. In other cases, clustering is performed beforehand during the analysis. In 10, the authors propose a clustering method for partitioning training paths based on key performance indicators (KPIs). Moreover, in 13, the authors segmented traces into three disjoint groups depending on the final grade mark and length of the study period. In both cases, the partitioning requires analyzing first such KPI and path lengths.
@@
When logs contain low-level information unsuitable for analyzing the problem, e.g., to analyze how knowledge areas spread over several courses are covered, it is possible to associate more precise semantics to those events (semantic lifting), e.g., associating the knowledge areas to the events 44. In 11, the authors propose a (semi)automatic procedure to associate semantics to training labels based on an ontology and then discover simplified training paths using the semantic information. This semantic lifting is also used in other works, such as 35, not to discover a trajectory but to perform a stopout/dropout analysis from another perspective.
@@
A curriculum process model can also be formally modeled from scratch. In 41, the authors propose a method based on Colored Petri Nets patterns and show how pattern mining can be used to discover a curriculum model.
@@
Curricular conformance checking is based on traditional conformance checking, except in 6, in which the authors check the curriculum design with the actual trajectories using sequence matching analysis. As described in 46, conformance is used in two senses: first, to identify how different cohorts behave and to identify outliers, and second, to measure adherence to a planned curriculum. As described in 19, trajectories have a high variance. Thus, two strategies are followed: clustering of traces once a cohort or specific group of students is identified, as described before, and projecting into the model and the log partial information or stages in the curricula, e.g., compulsory or first-year courses.
@@
In the case of bottleneck analysis, in all works 2, 9, 27, 30, a process model directly represents the students’ actual trajectories, i.e., each event is a course or exam taken or approved. Since time matters, it is firmly based on performance analysis. In 2, 27, the authors identify courses that prevent academic progress by performing a waiting time analysis and identifying courses that exhibit approval times above the expected. In 9, the authors obtained frequencies of students according to the time students invest in completing a course and extracted requirements that define what a bottleneck is from the observed behavior, e.g., courses with frequencies for total passed students lower than 50%. In 30, the authors detect exam-taking patterns and use prediction models to elucidate the impact of different factors on graduation times.
@@
In the case of stopout/dropout analysis, a process model can directly represent the students’ courses taken or approved. In 27, the authors follow this idea and identify the last courses with which students graduate or drop out. They consider if two years have passed since the last approval, a rule defined by their university. However, other works use different semantics for events. In 35, the authors built an event log at the semester level, converting approval data to higher-level events such as a STOPOUT event if the student did not enroll in courses in the semester and a BAG event if the student enrolls but has pending courses, among others. They also consider final events such as EARLY-DROPOUT, for students who dropped out no later than the fourth semester, and LATE-GRAD for students who took longer to graduate than the time defined for the program, among other events. In 36, the authors add other events of interest to see how pending courses accumulate in a semester and if that affects the stopout in the long run. They also studied how gender, income, and entry math skills may explain differences in trajectories. In 37, they refine this idea by defining a model according to the courses students have failed and have yet to pass. Finally, they focus on dropout in 38, using a state diagram that shows the possible result events in the retake-stopout model.
@@
Finally, trajectory/curriculum recommendations, such as in 29, 46, are based on combining several strategies. In general terms, as described in 46, PM techniques are used to identify successful students’ paths and to detect deviations from expected plans. This information is combined with rule-based AI strategies to derive rules as recommendations of more suitable study paths with higher success rates, i.e., predictions with a certain confidence level. Many possible strategies are applied, e.g., in 48, the authors apply sequence-based approaches (e.g., sequential pattern mining); in 25, the authors use logistic regression for resolving a classification problem; and in 32, the authors derive decision tree models. In a less restrictive curricular context, one possibility is to discover the trajectories that students naturally follow and define a curriculum based on that. In 10, the authors analyze social networks between courses and training providers of a consulting company, e.g., to discover popular providers. They also propose a clustering method for partitioning training paths based on key performance indicators and use discovery to analyze such clusters.
@@
In Table 4, there is a summary of the PM strategies followed, indicating the type of PM: Discovery (D), Conformance checking (C), and Extension (E). The primary type is discovery since most works tend to discover student trajectories (control-flow perspective). Few works also focus on the time perspective when considering enhancement 2, 29, 9, 27, e.g., bottleneck analysis, and only one paper 10 considers the organizational perspective to provide recommendations.
@@
As tangible results of the works, most of them provide real-world application examples by taking historical information from universities and consulting companies supporting software systems. From a methodological perspective, since 2019, some works have started using the PM2 43 methodology to guide the execution of their PM projects; before that, only a few works described using an ad hoc methodology.
@@
Only some works define comprehensive tools supporting curricular mining initiatives. In 29, the authors envision CurriM, a tool covering many PM actions about curricular models. Moreover, in 46, the authors describe an approach and tool to support individualized study planning and monitoring using campus management system data and study program models. In this case, the StudyBuddy tool provides a planning tool for students, while the BuddyAnalytics provides an analytics dashboard for program designers. Unfortunately, none of these tools are available. However, the project described in 46 is ongoing. Other works also refer to unavailable developed tools supporting their specific analysis.
@@
Before 2017, the most used tool for performing analysis was ProM. Some works describe the discovery algorithms used, i.e., Fuzzy Miner, 11, Heuristic Miner 11, 27 and infrequent Inductive Miner (iIM) 13. Moreover, in 4, the authors analyze the appropriateness of the Heuristic Miner for process discovery. Since 2017, there has been a slight change in the appearance of other PM tools, i.e., Disco and bupaR. Also, in 39, the authors describe the combination of PM and DM tools they used: KNIME, Disco, ProM, and R. Finally, in 47, the authors developed a curriculum simulator to mimic the behaviors of different kinds of students, which is unavailable.
Table 4: PM strategies.
Paper Year Result Tool Method PM Type Perspective
41 2009 example, tool ProM D, C control-flow
2 2012 example ProM Ad hoc D, C, E control-flow, time
4 2012 example ProM Ad hoc D control-flow
29 2012 tool ProM D, C, E control-flow, time
10 2014 example, tool ProM D, E control-flow, org.
11 2014 example ProM D, C control-flow
6 2015 example, tool ProM Ad hoc C control-flow
47 2015 tool D, C control-flow
13 2017 example ProM D, C control-flow
39 2017 example, tool ProM, Disco D control-flow
9 2018 example ProM Ad hoc D, E control-flow, time
48 2018 example, tool D control-flow
35 2019 example bupaR PM2 D control-flow
36 2019 example bupaR PM2 D control-flow
27 2021 example ProM D, E control-flow, time
37 2021 example bupaR PM2 D control-flow
38 2021 example bupaR PM2 D control-flow
25 2022 tool D control-flow
46 2022 example, tool D, C control-flow
30 2023 example Pm4Py D, C control-flow
19 2023 example ProM, Disco PM2 D, C control-flow
32 2023 example, tool Pm4Py D control-flow
4 Opportunities & challenges
@@
In this section, we discuss (RQ4) What are the open challenges and research opportunities? Since curricular analysis is within a broader context, there are general PM challenges from a technical 42 and organizational 26 perspective, as well as more specific EPM challenges 7 that must be addressed. In what follows, we summarize and complement specific curricular analysis challenges described in primary studies.
@@
PM tools must handle voluminous data from educational information systems. This data describes heterogeneous and complex situations since educational processes reflect the high diversity of behaviors in students’ learning paths 10. Additionally, some undesired behaviors from the PM perspective are not extraordinary, e.g., non-finalized cases of dropout students. Moreover, as described in 19, curricula are often flexible, particularly when there is a wide variety of compulsory courses and when exceptions are considered, e.g., exams could be repeated after failing, or there are lateral program entries when a student switches its career or spend a semester studying in a foreign program. All these aspects affect process discovery (generating spaghetti models) and conformance checking. Several strategies are followed, such as identifying consistent student cohorts or specific groups of students (clustering) with homogeneous characteristics and filtering them, reducing the complexity of the discovered process models.
Cluster analysis aims to identify typical behavior patterns of groups of similar students based on intra- and inter-course data such as examination results and order of the courses in their trajectories 24, 23. In 11, 13, the authors mentioned the need to develop new and refined clustering and classification techniques and to compare results from different techniques in the context of curricular analysis.
Clustering and filtering strategies could be applied on demand within an iterative analysis, e.g., performing trace clustering on students with similar characteristics and filtering a subset of such traces by looking at other specific student attributes. Another strategy is aggregating events when the raw data might need to be more cohesive or only considering partial information or stages in the curricula, e.g., compulsory or first-year courses, to perform the analysis. As noticed in 19, the use of filters to reduce the noise of the log and the resulting models comes at the cost of reducing their expressiveness.
@@
Some analyses, e.g., analyzing dropout, sometimes require links with other contextual aspects, such as demographic information of students and the context in which a course is offered (e.g., the number of teachers assigned to a course). In this context, two requirements arise: enriching logs with additional data and incorporating multidimensional and integrated analysis strategies between PM and data mining (DM), as devised in 14. DM allows projecting information from the observed data onto the models to understand underlying relationships 29. In this sense, it is possible not only to project simple demographic factors 17 such as gender, age, nationality, ethnicity, family education, or economic position but to complement and complete curricula analysis using the educational history of the students and vocational ambitions to explain differences in trajectories 34, and econometric models to quantify the influence of variables along time 35. DM can also enrich the analysis of university curricula and the acquired competencies, for example, adding Tracer Study (TS) data of their graduates, such as the waiting period for getting the first job, the success of graduates competing in the selection, earned salaries, and graduates’ opinions 3. Process Cubes 8 could be explored in this multidimensional context for data management and log generation.
Curricular flexibility and evolution
. Student diversity in higher education and the need to deliver students who adapt to changes has made educational institutions address these needs by designing a flexible curriculum 20. To answer if current prerequisites make sense, it is necessary to evaluate what would happen if these prerequisites were not present, something that can only be approximated based on assumptions or through the generation of simulations. Recent studies based on DM and simulation present generalized approaches to exploring the impact of policy changes on the curriculum and creating effective and individually verifiable simulations without requiring a rigid curriculum graph to evaluate changes to flexible curricula 5. Process simulation could be applied to possible trajectories based on the history of discovered trajectories. It would also be interesting to compare the results of both types of simulation to assess their potential.
Another related concept in the context of curricular evolution is concept drift 10, 29, 39, i.e., a significant change in a process over time. PM is particularly sensitive to this aspect, and more profound studies still need to be developed in the educational context.
@@
DM techniques have been widely used to predict students’ performance and dropout, using demographic and educational attributes, applying techniques such as Decision Trees, Random Forest, Bayes Network, and others 1. Since PM is also used for these purposes, comparing such results and determining how they can be combined is necessary.
@@
Using metrics provides objective information on which to evaluate the current state of a curriculum and base decisions regarding policy changes. It could be possible to combine the discovered trajectories with specific curriculum metrics 28. Only a few works, such as 46, have used Key Performance Indicators (KPIs) about students, study programs, or cohorts, such as the success rate of a course and exams passed in a semester. Further works could explore and evaluate the use of such information and determine its potential for integrated analysis.
@@
At a methodological level, some works use a well-defined but general methodology, e.g., PM2, for guiding a curricular analysis project. We may extend the PM2 process to conduct concrete curricular analysis activities. It would imply defining specific actions and recommending particular techniques to be applied in each step, according to predefined questions of interest for curricular analysis. Moreover, we can consider a methodological context with an analysis based on existing combined DM and PM methods, as in 14.
@@
Tools must be available if the curricular analysis is expected to be interpreted by the end users, e.g., CurriM 29 and StudyBuddy and BuddyAnalytics 46. Many opportunities exist, e.g., extending an existing PM tool with a specific layer of curricular analysis techniques guided by a concrete methodology.
Providing adequate visualization techniques and notations for improving the interpretation of results by the end users is also a challenge, as identified in 10, 7, 19.
There is also space for improvement in applying predictive PM to provide recommendations. It could be possible to provide students with the ability to simulate an ideal trajectory and predict the probability of success according to it.
Replication.
@@
Several aspects make replication of studies hard, e.g., no common methodology is followed for the same guiding questions, there are many alternatives to perform the analysis, data mining and PM strategies are usually disconnected, there is heterogeneous data, and no standard log format/content is defined, e.g., which demographic attributes should be considered for a given study. Some works, such as 46, tend to define more standardized aspects, e.g., a general data model for PM initiatives independent of academic systems and programs. Given that several referred study programs have similarities, based on a standard methodology and a set of predetermined studies, it would be interesting to perform a cross-university curricular analysis to detect similarities and differences.

### v2

@@ 38 D, C, P control-flow courses - -
As already described and summarized in Table 4, the main objective addressed in the papers is the discovery of educational trajectories. It requires defining a process model that reflects the trajectory of students. Cases are usually actual students’ trajectories in which events of the discovered process represent courses, e.g., in 54, 2, 4, 13, 45, 9, 30, 21, or certain curriculum milestones such as the end of a semester 33, 15, 16. However, the semantics of an event can be conceived from different perspectives, not only from a curricular standpoint, e.g., the courses a student approves, but also from a contextual perspective, e.g., graduation in time, late dropouts or the number of pending studies. This interpretation is used in works focusing on stopout/dropout analysis, e.g., 41, as we will describe later. Figure 4 depicts three different model notations with their own event semantics. The decision depends on different aspects. In 37, a curricular model represents courses (as transitions) with term start and end silent transitions, allowing the analysis of the temporal difference between the expected and actual course-taking term. In such a case, Workflow Nets are used since they allow the definition of a partial-order alignment algorithm using such notation formally. In 16, there is a representation of courses and milestones (end of semester and end of a year). However, in such a case, the authors chose Causal Nets since it provides a more compact and simpler representation of the routing logic inferred from the discovery algorithm (i.e., the Heuristic Miner). Finally, in 42, the authors model the context of a student as activities: enrollment, stopout/dropout, bag (i.e., pending failed courses), and empty (no pending courses). They use DFG since its notation is available in the statistical analysis environment they used.
((a)) Worflow Net from 37 (using Pm4Py)
((b)) Causal Net from 16 (using ProM)
((c)) DFG from 42 (using bupaR)
Figure 4: Example of process models for curricular analysis using different notations and event semantics.
@@
Since a trajectory refers both to a student’s path and a set of students, in some cases, there is an a priori clustering of traces based on considering additional information, e.g., demographic information or students’ performance characteristics, which sometimes requires additional data mining tools 41, 15. In other cases, clustering is performed beforehand during the analysis. In 10, the authors propose a clustering method, based on sequence clustering, for partitioning training paths based on key performance indicators (KPIs). Moreover, in 13, the authors segmented traces into three disjoint groups depending on the final grade mark and length of the study period, using trace clustering. In both cases, partitioning first requires analyzing KPI and path lengths.
@@
When logs contain low-level information unsuitable for analyzing the problem, e.g., to analyze how knowledge areas spread over several courses are covered, it is possible to associate more precise semantics to those events (semantic lifting), e.g., associating the knowledge areas to the events 51. In 11, the authors propose a (semi)automatic procedure to associate semantics to training labels based on an ontology and then discover simplified training paths using the semantic information. This semantic lifting is also used in other works, such as 41, not to discover a trajectory but to perform a stopout/dropout analysis from another perspective.
@@
A curriculum process model can also be formally modeled from scratch. In 47, the authors propose a method based on Colored Petri Nets patterns and show how pattern mining can be used to discover a curriculum model.
@@
Curricular conformance conformance checking. As described in 53, conformance is used in two senses: first, to identify how different cohorts behave and to identify outliers, and second, to measure adherence to a planned curriculum. As described in 21, trajectories have a high variance. Thus, two strategies are followed: clustering of traces once a cohort or specific group of students is identified, as described before, and projecting into the model and the log partial information or stages in the curricula, e.g., compulsory or first-year courses. The main conformance-checking technique used is alignment. However, there are a couple of exceptions. In 11, the authors propose using LTL expressions to check whether trajectories conform to specific requirements such as minimal paths and business rules. In 6, the authors check the curriculum design using sequence matching analysis to the actual trajectories using sequence matching analysis. Finally, in 15, the authors use a cost-based conformance-checking technique leveraging the fitness metric by considering the cost of skipping and inserting activities. An additional perspective to conformance-checking is followed in 37 in which partial orders are used to relax the constraints of strictly ordered traces, and alignments can be computed without requiring a strict order in which courses are offered.
@@
In the case of bottleneck analysis, in all works 2, 9, 30, 33, 16, 15, a process model directly represents the students’ actual trajectories, i.e., each event is a course or exam taken or approved. Since time matters, it is firmly based on performance analysis. In 2, 30, the authors identify courses that prevent academic progress by performing a waiting time analysis and identifying courses that exhibit approval times above the expected. In 9, the authors obtained frequencies of students according to the time students invest in completing a course and extracted requirements that define what a bottleneck is from the observed behavior, e.g., courses with frequencies for total passed students lower than 50%. In 33, 16, 15, the authors detect exam-taking patterns and use prediction models to elucidate the impact of different factors on graduation times.
@@
In the case of stopout/dropout analysis, a process model can directly represent the students’ courses taken or approved. In 30, the authors follow this idea and identify the last courses with which students graduate or drop out. They consider if two years have passed since the last approval, a rule defined by their university. However, other works use different semantics for events. In 41, the authors built an event log at the semester level, converting approval data to higher-level events such as a STOPOUT event if the student did not enroll in courses in the semester and a BAG event if the student enrolls but has pending courses, among others. They also consider final events such as EARLY-DROPOUT, for students who dropped out no later than the fourth semester, and LATE-GRAD for students who took longer to graduate than the time defined for the program, among other events. In 42, the authors add other events of interest to see how pending courses accumulate in a semester and if that affects the stopout in the long run. They also studied how gender, income, and entry math skills may explain differences in trajectories. In 43, they refine this idea by defining a model according to the courses students have failed and have yet to pass. Finally, they focus on dropout in 44, using a state diagram that shows the possible result events in the retake-stopout model.
Enhancement techniques usually overlay additional information in a process model to enhance comprehension of aspects. In 35, the authors introduce a method to associate different context representations of a data record to a label, i.e., the coloring of activities based on the case or event attributes and activity frequency. It differentiates many possible representations in the same data instance, e.g., color ranges represent the frequency of each course taken or the proximity to the minimum and maximum grades.
@@
Finally, trajectory/curriculum recommendations, such as in 32, 45, 53, 38, are based on combining several strategies. In general terms, as described in 53, 38, PM techniques are used to identify successful students’ paths and to detect deviations from expected plans. This information is combined with rule-based AI strategies to derive rules as recommendations of more suitable study paths with higher success rates, i.e., predictions with a certain confidence level. Many possible strategies are applied, e.g., in 55, the authors apply sequence-based approaches (e.g., sequential pattern mining); in 28, the authors use logistic regression for resolving a classification problem; and in 36, the authors derive decision tree models. In a less restrictive curricular context, one possibility is to discover the trajectories that students naturally follow and define a curriculum based on that. In 10, the authors analyze social networks between courses and training providers of a consulting company, e.g., to discover popular providers. They also propose a clustering method for partitioning training paths based on key performance indicators and use discovery to analyze such clusters.
@@
Table 7 provides a comparative view of process mining methods, tools, and algorithms used across the primary studies. Each entry includes the study’s title and the information explicitly provided concerning the methodology followed for the case study analysis (if applicable), the tool the authors developed (if applicable), and the base tools and discovery algorithms used.
Table 7: Process mining methods, algorithms, and tools.
Paper Method Tool Base tool Algorithm
47 StudyBuddy / BuddyAnalytics ProM, CPN Tools
2 Ad hoc ProM
4 Ad hoc ProM Heuristic Miner
32 CurriM ProM, CPN Tools
10 PHIDIAS ProM Fuzzy Miner, Social mining
11 ProM, WSML 2 Reasoner, LTL checker Heuristic Miner
6 Ad hoc Unnamed ProM
54
13 ProM Infrequent Inductive Miner
45 Unnamed ProM, Disco, KNIME, R, Shiny
9 Ad hoc ProM
55 Unnamed ProM
41 PM2 bupaR, R
42 PM2 bupaR, R
30 ProM Heuristic Miner, Inductive Miner
43 PM2 bupaR, R
44 PM2 bupaR, R
21 PM2 ProM, Disco
28 Unnamed
53 StudyBuddy / BuddyAnalytics
33 Pm4Py Infrequent Inductive Miner
36 AIStudyBuddy Pm4Py
15 Pm4Py Infrequent Inductive Miner
16 ProM Heuristic Miner
35
37 StudyBuddy / BuddyAnalytics Pm4Py
38 StudyBuddy / BuddyAnalytics
@@
As tangible results of the works, most of them provide real-world application examples by taking historical information from universities and consulting companies supporting software systems. From a methodological perspective, since 2019, some works have started using the PM2 50 methodology to guide the execution of their PM projects; before that, only a few works described using an ad hoc methodology. Although a methodology exists, the case studies do not always follow a systematic approach, and the methodology is not tailored for application in the specific context.
@@
Only some works define comprehensive tools supporting curricular mining initiatives. In 32, the authors envision CurriM, a tool covering many PM actions about curricular models. Moreover, in 53, 38, the authors describe an approach and tool to support individualized study planning and monitoring using campus management system data and study program models. In this case, the StudyBuddy tool provides a planning tool for students, while the BuddyAnalytics offers an analytics dashboard for program designers. Unfortunately, none of these tools are available. However, the project described in 53, 38 is ongoing. Other works also refer to unavailable developed tools supporting their specific analysis.
@@
Before 2017, the most used tool for performing analysis was ProM. Some works describe the discovery algorithms used, i.e., Fuzzy Miner, 11, Heuristic Miner 11, 30 and infrequent Inductive Miner (iIM) 13. Moreover, in 4, the authors analyze the appropriateness of the Heuristic Miner for process discovery. No other work compares algorithms to define the most suitable one in this context. The decision to use one or the other is usually not clearly justified, except in cases where the use of a specific notation is defined that favors some aspect, e.g., visualization of the results 16, or formal bases that allow defining additional algorithms 37. Since 2017, there has been a slight change in the existence of other PM tools, i.e., Disco and bupaR. Also, in 45, the authors describe the combination of PM and DM tools they used: KNIME, Disco, ProM, and R. Finally, in 54, the authors developed a curriculum simulator to mimic the behaviors of different kinds of students, which is unavailable.
4 Challenges & opportunities
@@
In this section, we discuss (RQ4) What are the open challenges and research opportunities? There are general PM challenges from a technical 48 and organizational 29 perspective, as well as more specific EPM challenges 7 that must be addressed. In what follows, we summarize and complement specific curricular analysis challenges described in primary studies.
Log abstraction
Log abstraction is a core preprocessing task in PM that simplifies logs by aggregating events, defining event classes, and clustering 27. In curricular analysis, however, process models must represent student trajectories at different levels of granularity, such as course-level activities versus curriculum milestones. An open challenge is to evaluate whether existing log abstraction techniques sufficiently capture these diverse granularity levels or if tailored, domain-specific methods are needed, as suggested by 35. Ensuring granularity and interpretability is critical to understanding student behavior across different levels.
@@
PM tools must handle voluminous data from educational information systems. This data describes heterogeneous and complex situations since educational processes reflect the high diversity of behaviors in students’ learning paths 10. Additionally, some undesired behaviors from the PM perspective are not extraordinary, e.g., non-finalized cases of dropout students. Moreover, as described in 21, curricula are often flexible, particularly when there is a wide variety of compulsory courses and when exceptions are considered, e.g., exams could be repeated after failing, or there are lateral program entries when a student switches its career or spend a semester studying in a foreign program. All these aspects affect process discovery (generating spaghetti models) and conformance checking. Addressing this challenge requires strategies such as clustering student cohorts with similar characteristics, using selective filtering to simplify process models without losing critical information, and developing more tailored conformance-checking techniques.
Cluster analysis aims to identify typical behavior patterns of groups of similar students based on intra- and inter-course data such as examination results and order of the courses in their trajectories 26, 25. In 11, 13, the authors mentioned the need to develop new and refined clustering and classification techniques and to compare results from different techniques in the context of curricular analysis. Clustering and filtering strategies could be applied on demand within an iterative analysis, e.g., performing trace clustering on students with similar characteristics and filtering a subset of such traces by looking at other specific student attributes. Another strategy is aggregating events when the raw data might need to be more cohesive or only consider partial information or stages in the curricula, e.g., compulsory or first-year courses, when performing the analysis. As noticed in 21, the use of filters to reduce the noise of the log and the resulting models comes at the cost of lowering their expressiveness.
New conformance-checking techniques could be analyzed, e.g., considering partial orders when computing alignments to make results less prone to the order in which courses are offered 37
@@
Some analyses, e.g., analyzing dropout, sometimes require links with other contextual aspects, such as demographic information of students and the context in which a course is offered (e.g., the number of teachers assigned to a course). Two challenging requirements arise: enriching logs with additional data and incorporating multidimensional and integrated analysis strategies between PM and data mining (DM), as devised in 14. DM allows projecting information from the observed data onto the models to understand underlying relationships 32. In this sense, it is possible not only to project simple demographic factors 19 such as gender, age, nationality, ethnicity, family education, or economic position but to complement and complete curricula analysis using the educational history of the students and vocational ambitions to explain differences in trajectories 40, and econometric models to quantify the influence of variables along time 41. DM can also enrich the analysis of university curricula and the acquired competencies, for example, adding Tracer Study (TS) data of their graduates, such as the waiting period for getting the first job, the success of graduates competing in the selection, earned salaries, and graduates’ opinions 3. Process Cubes 8 and object-centric process mining 49 could be explored in this multidimensional context for data management and log generation. Consider adding an analysis of external factors that may affect students’ educational trajectories could help distinguish between curricular and non-curricular reasons behind delays or dropouts.
Curricular flexibility and evolution.
Student diversity in higher education and the need to deliver students who adapt to changes has made educational institutions address these needs by designing a flexible curriculum 22. To answer if current prerequisites make sense, it is necessary to evaluate what would happen if these prerequisites were not present, something that can only be approximated based on assumptions or through the generation of simulations. Recent studies based on DM and simulation present generalized approaches to exploring the impact of policy changes on the curriculum and creating effective and individually verifiable simulations without requiring a rigid curriculum graph to evaluate changes to flexible curricula 5. Process simulation could be applied to possible trajectories based on the history of discovered trajectories. It would also be interesting to compare the results of both types of simulation to assess their potential.
Another related concept in the context of curricular evolution is concept drift 10, 32, 45, i.e., a significant change in a process over time. PM is particularly sensitive to this aspect, and more profound studies still need to be developed in the educational context.
@@
DM techniques have been widely used to predict students’ performance and dropout, using demographic and educational attributes, applying techniques such as Decision Trees, Random Forest, Bayes Network, and others 1. Integrating these predictive models with PM techniques is essential to leverage the strengths of both methods and improve prediction accuracy. Comparative analysis of PM and DM results is also needed to determine best practices for combining methods in educational prediction contexts.
@@
Using metrics provides objective information on which to evaluate the current state of a curriculum and base decisions regarding policy changes. It could be possible to combine the discovered trajectories with specific curriculum metrics 31. Only a few works, such as 53, have used Key Performance Indicators (KPIs) about students, study programs, or cohorts, such as the success rate of a course and exams passed in a semester. Further works could explore and evaluate the use of such information and determine its potential for integrated analysis.
@@
At a methodological level, some works use a well-defined but general methodology, e.g., PM2, for guiding a curricular analysis project. We may extend the PM2 process to conduct concrete curricular analysis activities. It would imply defining specific actions and recommending particular techniques to be applied in each step, according to predefined questions of interest for curricular analysis. Moreover, we can consider a methodological context with an analysis based on existing combined DM and PM methods, as in 14. Curricular analysis also requires an interdisciplinary approach that merges education, psychology, and data science insights, enabling more relevant and actionable results.
@@
Tools must be available if the curricular analysis is expected to be interpreted by non-technical users, e.g., CurriM 32 and StudyBuddy and BuddyAnalytics 53, 38. Many opportunities exist, e.g., providing real-time analysis tools and extending an existing PM tool with a specific layer of curricular analysis techniques guided by a concrete methodology. Providing adequate visualization techniques and notations for improving the interpretation of results by users is also a challenge, as identified in 10, 7, 21. There is also room for improvement in applying predictive PM to provide recommendations. Students could be able to simulate an ideal trajectory and predict the probability of success based on it.
Practical application.
@@
Several aspects make replication of studies hard, e.g., no standard methodology is followed for the same guiding questions, there are many alternatives to perform the analysis, data mining and PM strategies are usually disconnected, there is heterogeneous data, and no standard log format/content is defined, e.g., which demographic attributes should be considered for a given study. Some works, such as 53, describe more standardized aspects, e.g., a general data model for PM initiatives independent of academic systems and programs. Given that several referred study programs have similarities, based on a standard methodology and a set of predetermined studies, it would be interesting to perform a cross-university curricular analysis to detect similarities and differences.
Stakeholder feedback and validation are vital to ensuring the practical application of findings. This objective includes developing a feedback mechanism for sharing insights with curriculum designers, faculty, and advisors. These people can use the results to make informed decisions and validate the effectiveness of changes based on trajectory and curriculum recommendations.
Finally, all these analyses must be conducted strongly emphasizing ethical and privacy considerations. Ensuring data privacy and ethical use, particularly with sensitive student data, is critical. This objective outlines the methods for de-identifying data and securing permissions, supporting rigorous and ethical analysis.
Based on the current state of research and the open challenges we have described, we can identify several specific research opportunities for curricular mining summarized in Table 8. The main research directions include developing advanced data preprocessing methods tailored to educational data and addressing ethical considerations. Other directions focus on creating flexible, scalable, and interpretable models, integrating PM with predictive analytics to support at-risk students, and refining visualization tools in this specific domain (although much progress has been made in general-purpose mining tools). Developing adaptive recommendation systems and simulation techniques can also help model curricular changes. Cross-institutional analysis, metric development, stakeholder feedback, and concept drift analysis are essential for standardized methodologies and long-term adaptability, enabling more effective curriculum management and student support.
Table 8: Summary of research opportunities.
Category Research Opportunity Related Challenges
Data Preprocessing and Semantic Integration Develop tailored abstraction methods for curricular analysis. Heterogeneity and Complexity, Log abstraction
Integrate semantic context with models to improve interpretability.
Design clustering methods to group student patterns and focus on specific subgroups.
Addressing Ethics and Privacy Concerns Develop specific de-identification techniques and data security frameworks to ensure privacy compliance. Practical Application
Establish ethical guidelines for transparent and responsible use of data.
Flexible and Scalable Modeling Design hybrid models that balance flexibility and simplicity. Curricular Flexibility, Heterogeneity and Complexity
Develop adaptive models that automatically adjust to diverse educational contexts.
Predictive Modeling and Early Intervention Develop predictive models for dropout risks and academic support. Prediction
Incorporate detailed student data to improve accuracy in early warning systems.
Create early warning systems for timely dropout prevention.
Visualization and Bottleneck Analysis Tools Build real-time visualization tools to identify bottlenecks with interpretative insights. Heterogeneity and Complexity
Use ML to predict bottlenecks based on historical data.
Conformance Checking and Simulation for Flexible Pathways Develop metrics for varied student paths, focusing on sequence and performance. Curricular Flexibility and Evolution
Research hybrid approaches to balance adherence and outcomes.
Use simulations to model potential impacts of curriculum changes on pathways.
Standardized Methodologies, Tools and Stakeholders Feedback Create tools with visual and predictive capabilities accessible to non-technical users. Integrated Analysis Strategies, Methodologies, Practical Application, Tools
Develop interactive dashboards for student path exploration.
Include feedback systems for curriculum designers, faculty, and advisors.
Adapt PM2 methodology to guide educational analysis.
Develop validation methods and iterative feedback loops to ensure practical use of PM recommendations.
Explore interdisciplinary methodologies to enhance results.
Adaptive and Context-Aware Recommendations Research personalized recommendations based on student progression. Heterogeneity and Complexity, Integrated Analysis Strategies
Include demographic and external factors like financial or employment status to distinguish curricular impacts.
Develop multidimensional analysis approaches, e.g., exploring process cubes and object-centric process mining.
Metric Development and Success Evaluation Define outcome-based metrics (e.g., graduation rates) for curriculum effectiveness. Metrics
Measure the impact of curriculum changes on success metrics.
Concept Drift Analysis Create methods to detect changes in curricular processes over time. Curricular Flexibility and Evolution
Develop adaptive models that respond to ongoing curriculum updates.
Cross-Institutional interdisciplinary Analyses Standardize data models and methodologies to enable cross-university comparisons. Heterogeneity and Complexity, Methodologies, Practical Application
Develop a toolbox with basic techniques and guidelines for studying the main problems associated with curriculum analysis.